OpenStudy (anonymous):

Given the triangle and equation below, what is the value of x? sec∠C - cot∠A = 3x

OpenStudy (anonymous):

Without knowing the lengths of the sides of the triangle, this is difficult. However, a general case can be made.  Using these lengths specified, sec(C)-cot(A)=3x turns into: \[\sqrt{A^2+C^2}/C-A/C=3X\] \[\rightarrow \sqrt{A^2+C^2}-A=3X\] \[\rightarrow C^2=9X^2C^2+3AXC\] \[\rightarrow (9C)X^2+(3A)X-C=0\] Using the quadratic equaiton: \[\rightarrow X=(-A \pm \sqrt{A^2+4C^2})/(6C)\]